# Enhanced many-body quantum scars from the non-Hermitian Fock skin effect

### v1

@@ A.1 Effective non-Hermitian PXP model
Appendix B SII. Fock skin accumulation and hopping inhomogeneity in the FSA chain
@@
${\bf Primary~{}model:}\hat{H}_{b}=\sum_{j=2n}\hat{P}_{j-1}\hat{X}^{\prime}_{j}\hat{P}_{j+1}+\sum_{j=2n+1}\hat{P}_{j-1}\hat{X}_{j}\hat{P}_{j+1},$ (S13)
$\hat{H}_{c}=\sum_{j=2n}\hat{P}_{j-1}\hat{X}^{\prime}_{j}\hat{P}_{j+1}+\sum_{j=2n+1}\hat{P}_{j-1}(\hat{X}^{\prime}_{j})^{\dagger}\hat{P}_{j+1}.$ (S14)
@@
Figure S4: (a) Fidelity dynamics for the models in Eqs. (S12)-(S14), under the disorder of $W=1.8$ in Eq. (S17). We set $L=14$, $u=0.35$ and the |$\ket{\mathbb{Z}_{2}}$ initial state. Due to the absence of the Fock skin effect, the QMBS revivals for the model $a$ (blue curve) rapidly collapse. (b) The scaling behavior of the first fidelity peak $F_{1}$ within $t\in[2.5,7.5]$, obtained by $200$ realizations. We set $W=0.8$ and $u=0.4$ for (b).
@@
Appendix C SIII. Real-complex transition in the spectrum of the non-Hermitian PXP model
@@
Appendix D SIV. Entanglement entropy
@@
Appendix E Structure of scarred revivals
To further characterize the QMBS revivals discussed in the main text, we plot the following echoes in FIG. S8
@@ Appendix E Structure of scarred revivals
with $\bra{\psi(t)}=\exp(-i\hat{H}t)\ket{\psi(0)}$, states |$\ket{\phi}=\ket{\mathbb{Z}_{2}}$ and |$\ket{\phi}=\ket{\mathbb{Z}^{\prime}_{2}}=\prod_{i}\hat{X}_{i}\ket{\mathbb{Z}_{2}}$. For $u=0.0$, we find that many-body revivals act as an oscillation between $\ket{\mathbb{Z}}$ and |$\ket{\mathbb{Z}^{\prime}_{2}}$, with the fidelity decaying with each state in-step. Meanwhile, for the non-Hermitian case $u=0.2$, these oscillations become unbalanced, with echoes to |$\ket{\mathbb{Z}_{2}}$ being stronger than for the Hermitian $u=0.0$ case. This is further exemplified for $u=0.5$ and $u=0.8$, where the dynamics and revivals predominantly occur only for the |$\ket{\mathbb{Z}_{2}}$ state. This leads to the more complex oscillations observed in our model. Nonetheless, we emphasize that the lifetime of the echoes to the |$\ket{\mathbb{Z}_{2}}$ state are already greatly improved for $u=0.2$, where the dynamics still explore most of the Hilbert space. This means that the improvement is not merely due to the wave function localizing around the |$\ket{\mathbb{Z}_{2}}$ state.
@@
Figure S8: Echoes in Eq. (S25), with |$\ket{\phi}=\ket{\mathbb{Z}_{2}}$ for red curves and |$\ket{\phi}=\ket{\mathbb{Z}^{\prime}_{2}}=\prod_{i}\hat{X}_{i}\ket{\mathbb{Z}_{2}}$ for blue curves. For the Hermitian case with $u=0.0$, the dynamics exhibits balanced oscillation between states |$\ket{\mathbb{Z}_{2}}$ and |$\ket{\mathbb{Z}^{\prime}_{2}}$. Under the weak non-Hermiticity with $u=0.2$, the echo from |$\ket{\mathbb{Z}_{2}}$ state exhibits higher peaks, showing enhanced quantum scarring in our model. For the cases under strong non-Hermiticity under $u=0.5$ and $u=0.8$, more complicated oscillations emerge, and the echo of |$\ket{\mathbb{Z}^{\prime}_{2}}$ vanishes. System size is $L=14$ for all panels.

### v2

@@ A.2 Coherence time
Appendix C SIII. Fock skin accumulation and hopping inhomogeneity in the FSA chain
@@
${\bf Primary~{}model:}\hat{H}_{b}=\sum_{j\in{\rm even}}\hat{P}_{j-1}\hat{X}^{\prime}_{j}\hat{P}_{j+1}+\sum_{j\in{\rm odd}}\hat{P}_{j-1}\hat{X}_{j}\hat{P}_{j+1},$ (S22)
$\hat{H}_{c}=\sum_{j\in{\rm even}}\hat{P}_{j-1}\hat{X}^{\prime}_{j}\hat{P}_{j+1}+\sum_{j\in{\rm odd}}\hat{P}_{j-1}(\hat{X}^{\prime}_{j})^{\dagger}\hat{P}_{j+1}.$ (S23)
@@
Figure S6: (a) Fidelity dynamics for the models in Eqs. (S21)-(S23), under the disorder of $W=1.0$ in Eq. (S26). We set $L=14$, $u=0.35$ and the |$\ket{\mathbb{Z}_{2}}$ initial state. Due to the absence of the Fock skin effect, the QMBS revivals for the model $a$ (blue curve) rapidly collapse. (b) The scaling behavior of the first fidelity peak $F_{1}$ within $t\in[2.5,7.5]$, obtained by $200$ realizations. We set $W=0.8$ and $u=0.4$ for (b).
@@
Appendix D SIV. Real-complex transition in the spectrum of the non-Hermitian PXP model
@@
Appendix E SV. Entanglement entropy
@@
Finally, we consider the entanglement entropy under different values of $u$. As shown in FIG. S10, each eigenstate is identified through the overlap $\log(|\left\langle\psi|\mathbb{Z}_{2}\right\rangle|)$. It is evident that more low-entropy states emerge when $u>0.4$. When $u$ reaches $0.99$, these states accumulate near $S=0$. These results indicate a significant enhancement under large $u$, which is further supported by the results in FIG. S11.
@@
Figure S10: Half-chain entanglement entropy in Eq. (S22) under different values of $u$. The states with high $\mathbb{Z}_{2}$ overlap ($\log(|\left\langle\psi|\mathbb{Z}_{2}\right\rangle|)$) are colored in violet. Under the increasing values of $u$, the intrinsic scarred states show lower entropy, and there emerge more low-entropy states. The system size is $L=14$.
Appendix F Structure of scarred revivals
To further characterize the QMBS revivals discussed in the main text, we plot the following echoes in FIG. S11 (a)
@@ Appendix F Structure of scarred revivals
with $\bra{\psi(t)}=\exp(-i\hat{H}t)\ket{\psi(0)}$, states |$\ket{\phi}=\ket{\mathbb{Z}_{2}}$ and |$\ket{\phi}=\ket{\mathbb{Z}^{\prime}_{2}}=\prod_{i}\hat{X}_{i}\ket{\mathbb{Z}_{2}}$. For $u=0.0$, we find that many-body revivals act as an oscillation between $\ket{\mathbb{Z}}$ and |$\ket{\mathbb{Z}^{\prime}_{2}}$, with the fidelity decaying with each state in-step. Meanwhile, for the non-Hermitian case $u=0.2$, these oscillations become unbalanced, with echoes to |$\ket{\mathbb{Z}_{2}}$ being stronger than for the Hermitian $u=0.0$ case. This is further exemplified for $u=0.5$ and $u=0.8$, where the dynamics and revivals predominantly occur only for the |$\ket{\mathbb{Z}_{2}}$ state. This leads to the more complex oscillations observed in our model. Nonetheless, we emphasize that the lifetime of the echoes to the |$\ket{\mathbb{Z}_{2}}$ state are already greatly improved for $u=0.2$, where the dynamics still explore most of the Hilbert space. This can be extended to the long-time dynamics shown in FIG. S11 (b). These results indicate that the improvement is not merely due to the wave function localizing around the |$\ket{\mathbb{Z}_{2}}$ state.
As previously demonstrated in the main text and FIG. S11(a), there is a strong scar enhancement under large values of $u$, Building on this, we further investigated the dynamics in the presence of disorder $\sum_{j}w_{j}\hat{Z}_{j}$. Compared to the dynamics under $u=0.2$ [blue curve in FIG. S11 (b)], the green curve for $u=0.7$ in FIG. S11(c) shows robust revivals, while the red curve for $u=0.0$ decays rapidly. These results highlight the significant protection and stability of the enhanced scarred states.
@@
Figure S11: (a) Echoes in Eq. (S34), with |$\ket{\phi}=\ket{\mathbb{Z}_{2}}$ for red curves and |$\ket{\phi}=\ket{\mathbb{Z}^{\prime}_{2}}=\prod_{i}\hat{X}_{i}\ket{\mathbb{Z}_{2}}$ for blue curves. For the Hermitian case with $u=0.0$, the dynamics exhibits balanced oscillation between states |$\ket{\mathbb{Z}_{2}}$ and |$\ket{\mathbb{Z}^{\prime}_{2}}$. Under the weak non-Hermiticity with $u=0.2$, the echo from |$\ket{\mathbb{Z}_{2}}$ state exhibits higher peaks, showing enhanced quantum scarring in our model. For the cases under strong non-Hermiticity under $u=0.5$ and $u=0.8$, more complicated oscillations emerge, and the echo of |$\ket{\mathbb{Z}^{\prime}_{2}}$ vanishes. (b) Normalized fidelity $F(t)$ over long-time dynamics. Compared to the conventional scarring revivals (red curve), the non-Hermiticity of $u=0.2$ (dashed green curve) gives rise to robust enhanced revivals over an extended period of $t=120$. (c) Normalized fidelity $F(t)$ under the disorder $\sum_{j}w_{j}\hat{Z}_{j}$, with $w_{j}\in[-0.5,0.5]$. The green curve in (c) demonstrates the strong protection under disorders. System size is $L=14$ for all panels.
Figure S12: Left: The mean ratio of adjacent energy gaps $\left\langle r\right\rangle$. Right: The level spacing distribution. All data is for the fixed system size $L=28$. The red and violet dashed curves represent the the Gaussian orthogonal ensemble (GOE) and Gaussian unitary ensemble (GUE), respectively, while the black line is for the Poisson distribution of spacings, typical of integrable systems.
Appendix G SVI. Level spacing distribution
As a probe of ergodicity, we compute the average ratio of consecutive level spacings
$r_{n}={\rm Min}[\delta_{n},\delta_{n+1}]/{\rm Max}[\delta_{n},\delta_{n+1}],$ (S35)
where $\delta_{n}=E_{n+1}-E_{n}$ is the energy difference between two adjacent energy levels in ascending order. We recall that for $|u|<1$ the eigenvalues of the Hamiltonian are all real, therefore they can be sorted without ambiguity. To avoid degeneracies (or near degeneracies) between different symmetry sectors, we always restrict our study to the maximally symmetric sector. For $u=0$, this is the sector symmetric under translation and spatial inversion. For $u>0$, it is instead the sector symmetric under translation by two sites (due to odd and even terms of the Hamiltonian being different) and under the combination of spatial inversion and translation by one site. As shown in Fig. S12, while for $u=0$ the average gap ratio is not yet at the expected value for a chaotic system due to the slow convergence with system size of the PXP model [124], $\left\langle r\right\rangle$ quickly approaches $0.53$ upon increasing $u$. This matches the prediction for the Gaussian Orthogonal Ensemble (GOE). As expected, the gap ratio value then starts to dip again near the singular point $u=1$. We note that the dip of $\langle r\rangle$ shortly after $u=0$ is due to the translation symmetry being only weakly broken and is expected to disappear in the thermodynamic limit.
In addition to the level spacing ratio, we also show the full distribution of level spacings after unfolding for $u=0.8$ in Fig. S12, which again aligns well with the GOE expectation. In both panels of Fig. S12, we also show the prediction for an integrable system (Poisson) and for the Gaussian unitary ensemble (GUE), neither aligning with our data. Overall, the data is strongly suggestive that our model is robustly chaotic in the non-Hermitian regime as long as $u$ is not close to 1. The match with the GOE prediction also indicates that the Hermitian matrix with the same spectrum as our non-Hermitian one is invariant under time-reversal.